Q1. 3 moles of an ideal gas run through a Camot cycle (A-B-C-D-A) to carry out some work. Data is shown below in the table. Determine the followings. (Given: 7-1.8) (i) Unknown parameters (?) at each state Net work done in Joule with correct sign, Carnot efficiency and actual efficiency State Pressure (Pa) 1300000 Temperature (K) Volume (m) 0.011 0.027 0.018 A B C D

1
Expert's answer
2022-03-25T05:41:03-0400

Gas constant (R) = 8.314 mol/(J * K);

Isothermal expansion:

W = n * R * T * ln(VB/VA) = 3 * 8.314 * 298 * ln(0.027/0.011) = 7432.7 * ln(2.45) = 6674 J = 6.7 kJ;

W = n * R * T * ln(VC/VA) = 3 * 8.314 * 298 * ln(0.018/0.011) = 7432.7 * ln(1.63) = 3660 J = 3.7 kJ;

The Carnot efficiency describes the maximum thermal efficiency that a heat engine can achieve according to the second law of thermodynamics.

ηmax=1−TL/TH